OpenStudy (anonymous):

MEDAL AND FAN!!!!!!!!!!!!! Stephanie made 36 muffins. She made 24 apple muffins and the rest were banana. What is the ratio of apple muffins to banana muffins? A. 24 to 36 B. 2 to 1 C. 12 to 36 D. 2 to 3

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@rational @iGreen @Tootles143 and how did you get that answer

OpenStudy (anonymous):

36÷12=3 24÷12=2

OpenStudy (anonymous):

ok thx:)

OpenStudy (anonymous):

Np

OpenStudy (texaschic101):

NOT D she made 36 muffins in total........24 of them were apple...the rest (36 - 24 = 12) banana muffins made is 12 ratio of APPLE to BANANA is : 24 : 12 or 24/12 which reduces to 2/1 or 2:1

OpenStudy (anonymous):

ok?